DISPERSED FLUORESCENCE SPECTROSCOPY OF THE SiCN Ã ^{2}Δ – X̃ ^{2}Π TRANSITION

\underline{MASARU FUKUSHIMA} and TAKASHI ISHIWATA, Faculty of Information Sciences, Hiroshima City University, Asa-Minami, Hiroshima 731-3194, Japan.

We have generated SiCN in supersonic free jet expansions, and observed the laser induced fluorescence (LIF) of the Ã 2Δ – X̃ ^{2}Π transition. The LIF excitation and dispersed spectra have been reported previously^{a}. On this work, the LIF dispersed spectra from single vibronic levels were measured in higher resolution, and the vibrational structure of the X̃ 2Π state of SiCN was analyzed. Fermi resonances on the Si-CN stretching ν_{3} mode, n_{3} = 2 and 3, only at the Ω = \frac{3}{2} level have been observed. The Renner-Teller bending structure has been also analyzed.

aM. Fukushima and T. Ishiwata, 60th International Symposium on Molecular Spectroscopy, TC02 (2005).